# Bidirectional Task-Motion Planning Based on Hierarchical Reinforcement Learning for Strategic Confrontation

### v2

@@ V Conclusions
This paper has presented a bidirectional HRL approach to address the hybrid decision-making process in strategic confrontations. It has developed two-layer DRL networks that associate commands and actions with task allocation and path planning, respectively. Furthermore, the method has refined the interaction mechanism and introduced a cross-training method to enhance learning between the two layers. We also have incorporated a trajectory prediction model to convert abstract command representations into concrete action objectives. Experimental results have demonstrated that our method outperforms the baselines in terms of episode returns, decision time, and confrontation win rate. Finally, the generalization and adaptability of our method have been validated by applying the trained model to larger instances and real-robot systems, respectively.

### v3

@@ V Conclusions
This paper has presented a bidirectional HRL approach to address the hybrid decision–making process in strategic confrontations. It has developed two–layer DRL networks that associate commands and actions with task allocation and path planning, respectively. Furthermore, the method has refined the interaction mechanism and introduced a cross–training method to enhance learning between the two layers. We also have incorporated a trajectory prediction model to convert abstract command representations into concrete action objectives. Experimental results have demonstrated that our method outperforms the baselines in terms of episode returns, decision time, and confrontation win rate. Finally, the generalization and adaptability of our method have been validated by applying the trained model to larger instances and real–robot systems, respectively.